OpenStudy (anonymous):

Sketch the curve and find the arc length of the spiral r=e^(-theta) for theta >= 0

OpenStudy (amistre64):

the length of a curve is determined by adding up al the small little hypotenuses of right triangles along its length

OpenStudy (amistre64):

im sure there is a way to define in it polar tho but id have to lok for it :)

OpenStudy (anonymous):

i have the formula i kinda know how to do it here goes

OpenStudy (amistre64):

pg 580 has a nice little formula

OpenStudy (amistre64):

\[\int \sqrt{r^2+(r')^2}dt\]

OpenStudy (amistre64):

that latex is a little squished looking to me

OpenStudy (anonymous):

\[L=\int\limits_{a}^{b}\sqrt{r^2+r'^2}\]

OpenStudy (anonymous):

and i just plug it in i guess but i know dont know how to find the limits

OpenStudy (amistre64):

t >= 0 is al you get?

OpenStudy (anonymous):

my r is = \[e^-\theta\]

OpenStudy (amistre64):

you might have to do it in the same manner as an improper integral

OpenStudy (amistre64):

use t for theta, its just simpler :)

OpenStudy (anonymous):

yea thats all i get.. usually im given from 0<t<pi or somting

OpenStudy (amistre64):

e^{-2t} + {-e^t}^2 e^2t + e^2t = 2e^2t sqrt(2e^2t) = sqrt(2) e^t i think

OpenStudy (amistre64):

yeah; sqrt(e^2t) = e^t

OpenStudy (anonymous):

i got sqrt(2) (e^-t)

OpenStudy (amistre64):

yeah, that negative thing eluded me :)

OpenStudy (amistre64):

so lets try to integrate this thing from 0 to something, then see if we can take the limit of it as that something approaches infinity

OpenStudy (anonymous):

so my whole thing is \[\sqrt{2}[-e^-t]\]

OpenStudy (anonymous):

and evaluate next but i dont have the limits

OpenStudy (amistre64):

-sqrt(2) e^-t right -sqrt(2) e^n+sqrt(2) e^0 = -0+sqrt(2)

OpenStudy (amistre64):

the limits are from 0 to as this thing approaches infinity

OpenStudy (amistre64):

e^-inf = 0

OpenStudy (anonymous):

so will that be my answer? \[\sqrt{2}\]

OpenStudy (amistre64):

if thats all they give you is t >= 0 then yes

OpenStudy (anonymous):

ok cool thats all i see in the problems hahah

OpenStudy (amistre64):

\[\int_{0}^{inf}\]

OpenStudy (anonymous):

can i have a negative \[\sqrt{2}\]

OpenStudy (anonymous):

some how i got a -sqrt2